Status Report

NASA Space Station On-Orbit Status 12 Dec 2003

By SpaceRef Editor
December 12, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 12 Dec 2003

All ISS systems continue to function nominally, except as noted previously or below.

The crew was commended on yesterday’s Robotics activities. They completed more survey tasks than were expected. The parked Canadarm2 presently allows continued viewing of the ISS zenith side from the ground. [Activities will continue next week with an updated procedure that optimizes the SSRMS wrist joint cluster configuration for future external surveys, with the arm based at the MBS PDGF-1 (mobile base system–power & data grapple fixture #1). Next week’s activities will include the PDGF-1 walk-off (arm currently based at PDGF-2).]

The crew was given the Go for resuming walking/running workouts on the TVIS (treadmill with vibration isolation) at regular loads, after engineers reviewed the video and data from yesterday’s TVIS ACO (activation checkout). [The treadmill can be used in nonmotorized (passive) mode, with yaw & pitch axis stabilizers active and the roll-stabilizing gyro off, until March 1, 2004. During the last TVIS testing (12/4), structural dynamics loads on the Service Module (SM) were determined with IWIS (internal wireless instrumentation system) to be less than 5% of limit values; thus, TVIS/passive mode does not pose an immediate fatigue issue to ISS structure. An additional checkout with higher subject loading will be conducted in February, to test for the increased subject loading in March due to the crew’s heavier TVIS workout load starting 30 days prior to their return to Earth gravity.]

FE Alexander Kaleri performed routine maintenance on the SM toilet facilities (ASU), replacing their pretreat container (E-K) plus hose with a new assembly and discarding old one. Kaleri also recorded the counter (SPKU) reading of the toilet flush system (SP) before and after filling the E-K hose for downlink. [E-K contains five liters of pre-treat solution, i.e., a mix of H2SO4 (sulfuric acid), CrO3 (chromium oxide, for oxidation and purple color), and H2O (water). The pre-treat liquid is mixed with water in a dispenser (DKiV) and used for toilet flushing.]

Continuing the current round of monthly Russian Segment (RS) air ventilation systems preventive maintenance, Kaleri worked one hour in theFunktsionalnyi-Grusovoi Blok (FGB) for a cleanup of the protective mesh screens of its TsV ventilation fans. The fans were powered off for the task from the ground, later turned on again.

In preparation for next week’s R&R (removal & replacement) work on the RS SKV-2 air conditioner, which is to receive a new BTA heat exchanger, the crew today conducted a careful review of the planned activities, particularly the critical drainage and venting of the Freon-218 (Russian name: Khladon-218) coolant to vacuum on 12/15. [Khladon is a toxic Level 1 hazard. To protect against a leak/spill during the SKV-2 R&R, the crew will wear ZM half-masks with replaceable filter cartridges, currently stowed in the FGB.]

CDR/SO Michael Foale conducted the second 30-day functionality (“health”) test of the HRF GASMAP (Human Research Facility/Gas Analyzer System for Metabolic Analysis Physiology) of Increment 8. Last time done: 11/11. [After unstowing, cabling and powering up the equipment, Mike let it run for 2 hrs. in standby mode to warm up, then performed a health check and reconfigured the system for a 2-3hr. low power rundown prior to shutdown. Afterwards, GASMAP was deactivated, disconnected and stowed again. The consecutive real-time air sampling performed last time with PuFF software was not required this time.]

Mike used the CDMK (carbon dioxide monitoring kit) to take the weekly reading of the cabin air’s current CO2 partial pressure, for calldown to the ground (along with the battery status) for use in trending analyses. The CDMK was then restowed at the SM Central Post.

In preparation for the startup and operation of the U.S. CCAA (common cabin air assembly) air conditioner in the Airlock (A/L) on 12/15, to support the SKV-2 R&R, the crew removed stowed equipment from the A/L aft location to provide a 2-ft. clearance around the CCAA vents and smoke detector. [The bags and a short EMU were placed temporarily (for about a week) in the Node until further instructions.]

Sasha Kaleri prepared the daily IMS (inventory management system) “delta” file for automatic export/import to update the database; he also performed the daily routine maintenance of the SOZh life support system, while Mike Foale attended to the regular routine status checkup of autonomous Inc. 8 payloads.

In his weekly tagup with IMS specialists at TsUP/Moscow, the FE discussed open stowage issues, today focusing on the procedure for inventory management of Orlan spacesuit equipment, better control in the use of barcode labeling, and location/identification of specific items incorrectly listed in the database.

The crew removed the hard-disk drive (HDD) from the LAB RWS (robotics workstation) PCS laptop (which has a failed graphics card) and placed it into the A/L PCS shell that had a failed hard drive. The station is now again one PCS laptop above the Flight Rule limit for a minimum of two working PCSs. “Ghosting” of another HDD with a copy load is scheduled for 11/18.

The regular once-a-week maintenance reboot on the two other operational PCS laptops was performed by Kaleri, who also rebooted the OCA comm router laptop (every two weeks).

The crew worked out on TVIS, RED exerciser and VELO cycle with load trainer.

Mike transferred data files from the physical exercise equipment to the MEC (medical equipment computer) via memory card and RED log entries, for downlink on OCA comm. Later, he performed the periodic transfer of accumulated data files from the wristband HRM (heart rate monitor) receiver stations to the MEC for downlink, then deleting them on the HRM.

Kaleri completed another monthly sensor reading of the “Pille-MKS” radiation dosimeter, with its ten sensors placed at various locations in the RS (port cabin window, starboard cabin window, ASU toilet facility, control panel, etc.). Pille dosage values were to be called down to MCC-M or downlinked via Regul comm. [Last time done: 11/13.]

The Science Officer completed the microbial analysis of water samples collected last week from the SM potable water ports with the U.S. WS&A (water sampler & archiver) and the WMK (water microbiology kit) at the T+5d incubation point and then loaded the data into the MEC.

At 9:00am EST, the crew set up and conducted a 10-min. ham radio session with amateur radio fans at theBerufliche Schule des Kreises Nordfriesland (Vocational School of Province Nordfriesland) in Niebull, in the most northwestern part of Germany. [Questions by students were uplinked beforehand. The Berufliche Schule in Niebull has 2600 students and 135 teachers.]

Other opportunities for ham radio passes have been identified for tomorrow and Sunday (12/14) for Australia/New Zealand and Japan, and provided to the crew. [Amateur radio activities are spearheaded by an organization formed by national and international ham radio groups called ARISS (Amateur Radio International Space Station). Russia has provided ports on the SM for radio antennas and the new Kenwood “Sputnik” system, and ISS crews have trained to operate the equipment. The first initial radio station was flown on STS-106/2A.2b in September 2000 and transferred to the ISS. ISS operations at present use voice and Packet, the Russian text-messaging device.]

At 10:05am, Foale and Kaleri held a lively PAO/educational TV exchange with attendees at a “Centennial of Flight” Educational Event at the Wright Brothers National Memorial, Kill Devil Hills, NC, talking with students from Kill Devil Hills and Kitty Hawk schools and displaying a model of the Wright Flyer (floating in zero-G), built by middle school students from Orono, Maine. [The theme of the day’s events, sponsored by NASA and the National Parks Service, is “Igniting the Imagination”, and the event is part of the Centennial of Flight (CoF) celebration at the Wright Brothers National Memorial, Dec. 12-17. CoF is a national celebration commemorating the first powered, heavier-than-air flight by the Wright brothers on Dec. 17, 1903.]

The S1 truss segment’s ETVCG (external TV camera group) camera sustained another loss of video today. There have been several such occurrences over the past week, all of which were successfully recovered with a power cycle. This anomaly continues to be investigated.

CMG Watch: CMG (control moment gyroscope) operations continue to be closely monitored. This morning (~3:30am EST), CMG3 spin motor current increased to 0.409 amp when the outer gimbal’s angle moved from zero degrees to >45 deg, then stabilized at 0.43 amps. This signature is similar to what was observed earlier this week. It is expected that when the outer gimbals angle returns to near zero the current will return to 0.36A. Evaluation continues. [GN&C (guidance, navigation & control) specialists have established specific constraints for CMG temperatures and spin motor currents as well as guidelines for mitigating risk and prolonging the life of the gyros. Anytime the identified signatures occur, high-rate data dumps will be performed to computer storage, then downlinked. As an example, if accelerometer output from a CMG reaches 0.2 g, the CMG will be taken out of the active steering loop, to ensure zero gimbal rates. If, in addition, the current drawn by its spin motor is 0.5 amps or greater for five minutes, the rotor will be commanded off. This guideline is derived from the experience with the failed CMG-1, and there is a list of other precautionary “if-then” criteria.]

Consumables Watch: Onboard consumables remain in good shape, generally tracking predictions well. [Water use remains slightly above the 2 liters/day projection, and six EDVs (132 liters) with H2O will be manifested on Progress 13P. Food, both U.S. and Russian, will also be shipped up on 13P (with 13P rations, US food is projected to last until March 20, 2004). Propellant supplies are more than adequate, particularly since an originally planned reboost in October was not performed (13P will bring 1300 kg props). 13P manifesting is also being conducted for two advanced SchRED (Schwinn resistive exercise device) canisters capable of 75000 cycles, as well as for two TVIS gyro bearings and other hardware, a new MCA (major constituents analyzer), etc.]

Procedures for best utilization of the 21 kg (12 days) of oxygen (O2) supply stored in Progress 12P have been worked out with Moscow in a 12P O2 Repress Plan. [Current SM GA (gas analyzer) ppO2 reading is 150.4 mmHg; the U.S. CSA-CP (compound specific analyzer-combustion products) indicated 157.5 mmHg at last reading. These O2 partial pressure levels are sufficient for crew health. 12P represses (or Elektron production increase) are not planned until GA or CSA-CP shows ppO2 readings below 146 mmHg (allowable minimum: 120 mmHg). Next 12P repress is scheduled for next week (12/17), with MCA to be activated ~12 hours before and calibrated ~6 hours before the repress. Other steps follow; to be repeated until all 12P O2 is used up.]

Today’s CEO targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, were Mt. Cameroon, Cameroon (Dynamic event: Equatorial cloud has migrated southward so that ISS/CEO may have been be able to image lava flows on this major active volcano that is seldom photographed. Looking just left for this high mountain that slopes directly into the sea. Mt. Cameroon is part of a chain of volcanoes that extends SW as a series of islands), Bamako, Mali (nadir pass over Mali’s capital city, just upstream of the inland delta of the Niger River, and a center of Sahelian and cross-Saharan trade for centuries), Tuamotu-Austral Islands (looking near nadir for atolls with coral reefs), El Paso, Texas (nadir pass. Trying to include the larger Mexican half of the urban region on the S side of the Rio Grande), Major snow streak, Midwest (Dynamic event: Thick snow has fallen in a wide swath from Denver to the Upper Peninsula of Michigan), Melbourne, Australia (orbital precession and sleep times have locked Australia out of targets lists for months, except for occasional crossings. There is still enough summer illumination for a shot of Melbourne), SE Australian floods (Dynamic event, using 400-mm lens. Suggested was a nadir mapping swath to document widespread flooding that made world headlines), Sydney, Australia (nadir pass), Johnston Island Reef, Pacific (looking a touch left. Coral reefs are the subject of interest on this 1000 yard-long atoll. During the 1950s and 1960s, the US Air Force conducted a dozen nuclear-test launchings. Two missiles exploded over the runway. Since then, the US collected 60, 000 cubic yards of radioactive material. The Fish and Wildlife Service is scheduled to gain oversight of the island in 2004 as a wildlife refuge), Nihoa reef, Pacific (nadir pass), and Lake Eyre, Australia (water levels in Lake Eyre have not been documented by handheld camera for many months).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 1:20pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On, 20A. Vozdukh CO2 scrubber is On (in Manual Mode 5/3). U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode). TCCS (trace contaminant control subsystem) is operating. MCA (major constituents analyzer) is off (in Life Extending Mode). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment: Pressure (mmHg) — 742; temperature (deg C) — 25.5; ppO2 (mmHg) — 149.0; ppCO2 (mmHg) — 3.9.
  • SM Transfer Compartment: Pressure (mmHg) — 758; temperature (deg C) — 19.8.
  • FGB Cabin: Pressure (mmHg) — 756; temperature (deg C) — 21.7.
  • Node: Pressure (mmHg) — 747.71; temperature (deg C) — 23.3 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab: Pressure (mmHg) — 749.97; temperature (deg C) — 24.7; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock): Pressure (mmHg) — 750.07; temperature (deg C) — 26.8; shell heater temp (deg C) — 24.5, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1: Shell heater temp (deg C) — 22.2
  • PMA-2: Shell heater temp (deg C) — 10.6

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in Directed Position (non-suntracking, “night glider”/”sun slicer” drag reduction mode).
  • SM batteries: Battery #8, formerly known as #7, is still disconnected in slot #8 for troubleshooting, off-line; all other batteries (7) are in “Partial Charge” mode.
  • FGB batteries: Battery #5 is off (capacity restoration mode, ROM); battery #1 is in “Cycle Mode”; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available:3680 kg (8113 lb) as of 12/11/03 [SM (755) + FGB (2573) + Progress M (352) + Progress M-1 (0)]. (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: -89.9 deg, pitch: -8.9 deg, roll: 1.8 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #2/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings.
  • MT: latched and mated at WS4.
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit (as of this morning, 9:03am EST [= epoch]):

  • Mean altitude — 371.7 km
  • Apogee — 375.7 km
  • Perigee — 367.7 km
  • Period — 92.0 min.
  • Inclination (to Equator) — 51.63 deg
  • Eccentricity — 0.0005921
  • Orbits per 24-hr. day — 15.65
  • Solar Beta angle: -59.6 deg (magnitude increasing).
  • Mean altitude loss in last 24 hours — 55 m
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 28889
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.